# Vid2Robot: End-to-end Video-conditioned Policy Learning with Cross-Attention Transformers

### v1

@@ Abstract
While large-scale robotic systems typically rely on textual instructions for tasks, this work explores a different approach: can robots infer the task directly from observing humans? This shift necessitates the robot’s ability to decode human intent and translate it into executable actions within its own physical constraints and environment.
We introduce Vid2Robot, a novel end-to-end video-based learning framework for robots. Given a video demonstration of a manipulation task and current visual observations, Vid2Robot directly produces robot actions. This is achieved through a unified representation model trained on a large dataset of human video and robot trajectory. The model leverages cross-attention mechanisms to fuse prompt video features to the robot’s current state and generate appropriate actions that mimic the observed task. To further improve policy performance, we propose auxiliary contrastive losses that enhance the alignment between human and robot video representations.
We evaluate Vid2Robot on real-world robots, demonstrating a 20% improvement in performance compared to other video-conditioned policies when using human demonstration videos. Additionally, our model exhibits emergent capabilities, such as successfully transferring observed motions from one object to another, and long-horizon composition, thus showcasing its potential for real-world applications. Project website: vid2robot.github.io.
@@ I Introduction
The path to creating versatile robots that provide assistance in people’s daily routines requires them to learn new skills on-the-go. These range from small preferences like which brand of dishwasher a specific household uses to entirely different ways to clean the house. For known skills, humans simply communicate in natural language, but when nuance is required or a skill is novel, we revert to demonstrations. For example, we might show how a particular microwave works or how we prefer our cabinets to be organized. To enable seamless robot deployment, robots need the same ability for generalization from demonstration for learning new policies that comes so naturally to humans.
@@
Humans can infer the intentions of other humans based on third-person visual observations. Oftentimes, we use social reasoning and common sense to understand others’ goals implicitly. This ability is leveraged both as children and adults (e.g. via How-To videos [30]) for learning anything where the mechanical nuance of the task is hard to capture in still images or text [5] (e.g. how to knead dough or knit). If robots can also be taught to understand the intentions of other agents, it might allow them to better interact with humans and perform tasks more efficiently.
@@
This work focuses on visual imitation learning, where robots learn to perform tasks by watching video demonstrations. This setup offers several advantages. First, it allows robots to learn from agents with different embodiment, enabling new skill acquisition without tele-operation. Second, it enables robots to learn from experts, even if they are not situated with the robot. Finally, visual imitation learning is ideal for teaching tasks that are difficult or impossible to describe in words.
@@
Existing multi-task robot manipulation models (e.g. RT-1 [7], RT-2 [8], and RT-X [33]) use language conditioning to output a robot trajectory. This reliance of text alone for task specification makes it difficult for robots to handle polysemy and tasks whose executions vary dramatically based on context. For example, ‘open drawer’, ‘open cabinet’, ‘open container with lid’ and ‘open jar with screw cap’ might share the same verb but require very different motor control for each interaction. Here the agent should not generalize its policy, whereas it should generalize from one drawer to others that vary in type, color and shape. For this reason, there are a broad range of tasks for which it is hard to design primitives for high-level planning approaches [25, 2].
Another common approach has been to use a final goal image in goal-conditioned behavior cloning tasks [31, 24]. While several task specifications can be defined in terms of the resulting state of the environment, there are others for which the manner in which the action is performed is part of the specification. For example, ‘hold the flag’ and ‘wave the flag’ can have the same final goal image. This ambiguity can be resolved through the use of several sub-goal frames, that is, video conditioning.
While language conditioned policies achieve somewhat high success rates, video conditioned policies have lagged behind in performance, as shown in prior work [21]. Cases of good performance [40] with video conditioning require the provided video to be from the same workspace with limited variability. Based on observations, we identify three main challenges for video conditioned policies: (1) High dimensional data: Raw videos are high dimensional data that require more compute and memory to process. This makes video conditioned multi-task policies difficult to train at scale. (2) Variability in task specification: There can be significant variance in how people perform the same task. Demonstrations for a task like ‘unstack the cups’ can have both visually distinctive and physically diverse cups, in addition to changes in the background distractors and lighting conditions. This leads to high variability in task specification for a policy that needs to perform the same task in a new setting. (3) Limited Availability of Training Data: While there is an abundance of unlabeled video data on the internet, obtaining labeled video datasets for specific tasks that our robots are capable of doing is challenging.
Despite these challenges, as noted, video conditioned policy learning is a core challenge robots need to master. Therefore, to reduce the reliance on detailed and potentially ambiguous language prompts, we aim to enable physical visual demonstrations as a another way for task specification. To this end, we study how end-to-end models with video-conditioning can used to specify tasks to robot.
@@
We aim to develop an end-to-end system that enables rapid adaptation to tasks specified in the form of video demonstration. Unlike prior work that either learned representations from videos for only object and verb recognition [21] or learned motor control in simulation [44], our work demonstrates the applicability of end-to-end learned video representations for real-world multi-task robotic control. We present the key contributions of our work as follows: (1) We present a transformer-based policy to encode video task specification, demonstrated by either robot or human agent embodiments (§II). (2) We encourage alignment between the prompt and robot video representations using three contrastive auxiliary losses during training (§II-E) (3) Through real robot experiments, we find our video conditioned policy is better than baselines on human prompt videos. Furthermore, our policy is better at cross-object motion transfer (§III).
@@
Figure 2: Dataset creation. (top row) Here we show a Robot-Robot video pair for placing the rxbar into top drawer. We similarly pair existing robot-robot videos performing the same task. (middle row) Here we show Hindsight Human-Robot paired videos for picking a coke can from the bottom drawer and placing it on the counter task. We use the task instructions from robot trajectories and ask human participants to perform the task and record a demonstration video from robot’s perspective/view. (bottom row) Here we show a Co-located Human-Robot pair of videos for placing the pipe wrench in the toolkit. We record both a human demonstration and by a robot teleoperation in a same workspace. Different workspaces can be used to perform the same task instruction, thus, eventually resulting in pairs with visually diverse prompts and robot state observations. More details in §II-B.
@@ II-A Preliminaries
Our objective is to design a robotic system that takes in a prompt video of a manipulation task and outputs actions that accomplish the task demonstrated in the video. This system needs to infer the underlying task from the prompt video (which might have a different setup or embodiment than the robot) and then manipulate the objects in its own environment to achieve the inferred task. Specifically, we are given a prompt video $V$ and the robot state $S_{t}=\{x_{i}\}_{i=t-k-1}^{t}$ where $x_{i}$ is the frame from the robot’s camera stream at time $i$, $k$ is the maximum number of historical frames, and $t$ is the current time-step. We train a policy $\pi(a_{t}|S_{t},V)$ that infers the underlying task from $V$ and predicts task relevant action $a_{t}$. To train this model, we need a dataset of paired prompt videos and robot trajectories. We will discuss in detail how to create paired datasets below.
@@ II-B Datasets
To train a video-conditioned robot policy we need a dataset of pairs: prompt videos and robot trajectories performing the same task. In this work, we explore prompt videos where the task is performed by both humans and robots. To create this dataset, we rely on three classes of data:
  1. 1.
@@
    Robot-Robot: We pair existing robot-robot videos of the same task. For this pairing we consider two videos to match if they are performing the same task in different settings. We define ‘task‘ based on natural language instructions used when recording robot trajectories. These instructions typically consist of one or two verbs surrounded by nouns, such as ‘place water bottle upright’, ‘move the coke can to the green chip bag’ or ‘open top drawer’. The objective of this pairing is two-fold: first, to be able to take advantage of an already labeled and collected dataset of robot trajectories and second to ensure robots are able to imitate when the same task is demonstrated in a different environment.
  2. 2.
@@
    Hindsight Human-Robot: Here we use the task instructions from the robot trajectories dataset and ask one to five human participants to perform the task and record a demonstration video from the robot’s perspective/view. The set of instructions are the same as before, but there is a significant embodiment and speed variability due to different humans performing the task with left or right hands and at a randomized robot camera angle. This requires some manual effort but provides us with a lot of paired data for training the policy for the available set of instructions in the robot dataset without having to collect new robot trajectories.
  3. 3.
@@
    Co-located Human-Robot In this case, a human and a robot perform the same task in the same workspace. We used this approach to collect human demonstrations and robot trajectories in diverse spaces such as a living space with sofas, a meeting room with whiteboards, hardware workstations with toy tools, a kitchen with a countertop, refrigerator and sink, a storage supplies area, and more.
@@
We show examples of paired prompt and robot videos from each of the three datasets in Figure 2. As can be seen, there is a considerable difference in the backgrounds and distractor objects in the Hindsight Human-Robot and Co-located Human-Robot datasets. A different complexity arises when comparing the first approach (Robot-Robot) where the actor is a robot with same morphology to the other two cases where the human is the actor in the prompt videos.
After combining all the datasets, we have $\sim$100k robot videos and $\sim$10k human videos covering the tasks introduced in RT-1 [7] and RT-2 [8]. We include videos from all three data sources as they represent varying levels of difficulty and expense to collect. Pairing existing robot datasets requires less additional effort but lacks diversity in how the task is done. The second source of data is created by asking humans to mimic existing robot trajectories. While this adds some diversity in prompt videos, it does not cover any new tasks on the robot side. Finally, the presumed gold-standard is to collect data where both humans and robots are co-located in the same environment and perform diverse tasks. This takes the most amount of time as labor is required both of the humans and robot trajectories collected through tele-operation.
@@ II-C Model Architecture
Figure 3: Architecture. Our model takes as input frames of the prompt video and the robot’s current observations, encodes those into prompt video and robot state token embeddings, which are then processed through into state-prompt encoder and decoded into a robot action for the current timestep. More details in §II-C.
@@
Our policy takes as input the prompt video and the current robot state and outputs robot actions. It consists of four modules: (1) prompt video encoder (2) robot state encoder, (3) state-prompt encoder, and (4) robot action decoder. The full architecture is illustrated in Figure 3 and each of the modules are detailed below:
@@
(1) Prompt Video Encoder encodes the video demonstration provided as a reference to convey the desired task semantics. The prompt video encoder implicitly learns to infer what task should be performed and how it needs to be done. The prompt encoder consists of a per-frame Image encoder $\phi_{p}$ (ViT [14]) followed by a Perceiver Resampler [1, 19] $\psi_{p}$. The output of the prompt encoder $\psi_{p}(\phi_{p}(V))=z_{prompt}$ is a set of $N$ tokens of d-dimension to condition the policy with the task relevant attributes from the video.
@@
(2) Robot State Encoder encodes the current state of the robot given the current frame and last $k$ frames as input. Note that this module also encodes information about the objects and environment of the robot. The architecture is similar to the prompt encoder, that is, a per-frame Image encoder $\phi_{s}$ followed by a Perceiver Resampler $\psi_{s}$. Similar to the prompt encoder’s outputs, the output of the state encoder is $\psi_{s}(\phi_{s}(S_{t}))=z_{state}$ that encodes the latent environment and robot state information from the history of recent observations.
We use the same image encoder weights for both (1) and (2), that is, $\phi_{p}\!=\!\phi_{s}\!=\!\phi$. The role of the image encoder $\phi$ is to capture spatial visual information in each frame. The Perceiver Resampler is used to enable temporal learning across frames as well as reduce the number of video tokens that must be passed into the action decoder.
@@
(3) State-Prompt Encoder The state-prompt encoder takes the prompt video encoding $z_{prompt}$ and robot state encoding $z_{state}$ and outputs a task encoding relevant for action prediction $z_{state|prompt}$. The module is trained to output robot actions by cross-attending between the state encoding as queries and the prompt video encoding as keys and values. Intuitively, the state-prompt encoder enables fusion of the state and prompt information. For example, if the prompt video demonstrates picking up of an apple in the basket and the current state contains apple, banana and orange, then the cross attention between the state and prompt encoding enables learning for which object to attend to in the state, which is crucial for the next step of action decoding. We refer to the output of the state-prompt encoder as prompt-aware state tokens.
@@
(4) Robot Action Decoder The goal of the action decoder is to predict the action vector $a_{t}$ for the current state $S_{t}$ such that it completes the task shown in the prompt video $V_{p}$. The action decoder is a transformer decoder architecture that takes in the fixed action position tokens [49] as input queries and the prompt-aware state tokens $z_{state|prompt}$ for keys and values. The size of the action position embedding is $N\times d$ where $N$ is the number of action dimensions and $d$ is the transformer embedding dimension. More details on the action vector in §II-D.
@@
The action position embeddings cross-attend to the prompt-aware state tokens to predict the target binned action values as output. Each output token of the action decoder corresponds to an action dimension for the mode, arm and base. Specifically, each token embedding is projected to 256 dimensions and a softmax layer is applied on the top to obtain the bin corresponding to the target action vector. Unlike prior work [7, 8] that use autoregressive action decoding that requires multiple forward passes during inference, we use action position embeddings for one forward pass prediction like in ACT [49]. Instead of predicting one action for the next timestep, we follow the approach outlined in [21, 49] and train the policy with a prediction horizon of four steps. We always use the action bin that has the highest probability, i.e. argmax over predicted probabilities, to choose the action value for execution.
Cross-Attention Layers. In the Vid2Robot architecture, we use Cross-Attention Transformer layers extensively. They are used in the following modules: Prompt Resampler, State Resampler, State-Prompt Encoder and Action Decoder. We found Cross-Attention layers are helpful in managing the high number of tokens and the resulting large attention matrices when processing both prompt videos and robot state videos. This is because the standard self-attention layers would require orders of magnitude more memory to process the same video. For example, when using ViT-B/16 the total number of video tokens for a $16$ frame reference video and a $8$ frame robot state video at $224\times 224$ resolution would be $8\times 196+16\times 196=4704$. A full self-attention operation on this would lead to an attention matrix with $4704^{2}\sim 22\mathrm{M}$ entries. However, by using two Perceiver Resamplers with 64 latents we were able to train with attention matrices of the size $8\times 196\times 64+16\times 196\times 64\sim.3\mathrm{M}$. Thus, cross attention layers in Vid2Robot play an important role in reducing attention computation and enabling training with paired videos.

### v2

@@ Abstract
Large-scale multi-task robotic manipulation systems often rely on text to specify the task. In this work, we explore whether a robot can learn by observing humans. To do so, the robot must understand a person’s intent and perform the inferred task despite differences in the embodiments and environments. We introduce Vid2Robot, an end-to-end video-conditioned policy that takes human videos demonstrating manipulation tasks as input and produces robot actions. Our model is trained with a large dataset of prompt video-robot trajectory pairs to learn unified representations of human and robot actions from videos. Vid2Robot uses cross-attention transformer layers between video features and the current robot state to produce the actions and perform the same task as shown in the video. We use auxiliary contrastive losses to align the prompt and robot video representations for better policies. We evaluate Vid2Robot on real-world robots and observe over 20% improvement over BC-Z when using human prompt videos. Further, we also show cross-object motion transfer ability that enables video-conditioned policies to transfer a motion observed on one object in the prompt video to another object in the robot’s own environment. Videos available at vid2robot.github.io.
@@ I Introduction
The path to creating versatile robots that assist in people’s daily routines requires them to learn new skills on-the-go. These skills can vary from simple preferences for arranging dishes in the dishwasher in a specific household to completely different approaches to household cleaning. Humans can communicate in natural language for tasks that are already known. However, we revert to demonstrations when we want to learn a novel skill with nuance. For example, we might show how a particular microwave works or how to organize our cabinets. Robots need the same ability for generalization from demonstration, which comes so naturally to humans.
@@
Humans can infer the intentions of other humans based on third-person visual observations. Often, we use social reasoning and common sense to understand others’ goals implicitly. This ability is crucial for learning everyday tasks, such as kneading dough or knitting, where the intricacies are challenging to convey through still images or text [6]. We often turn to How-To videos [32]) to learn how to perform such tasks. If robots could act based on videos, it would enable efficient task learning and improved interaction with humans.
@@
This work focuses on visual imitation learning, where robots learn to perform tasks by watching video demonstrations. This setup offers several advantages. First, it allows robots to learn from agents with a different embodiment, enabling new skill acquisition without teleoperation. Second, it allows robots to infer tasks from expert demonstrations, even if the expert is not showing how to perform tasks in the same environment as the robot. Finally, visual imitation learning is ideal for teaching tasks that are difficult or impossible to describe in words.
@@
Existing multi-task robot manipulation models (e.g. RT-1 [8], RT-2 [9], and RT-X [35]) use language conditioning to output a robot trajectory. Relying on text alone for task specification makes it difficult for robots to handle polysemy and tasks whose executions vary dramatically based on context. For example, ‘open drawer’, ‘open cabinet’, ‘open container with lid’ and ‘open jar with screw cap’ might share the same verb but require very different motor control for each interaction. Here, the agent should not generalize its policy, whereas it should generalize from one drawer to others that vary in type, color, and shape. For this reason, there are a broad range of tasks for which it is hard to design primitives for high-level planning approaches [27, 2].
[…]
Despite these challenges, video-conditioned policy learning is a core challenge robots need to master. To this end, we study how end-to-end models with video-conditioning can be used to specify tasks to the robot. Vid2Robot is an end-to-end system that enables rapid adaptation to tasks specified as video demonstrations. Unlike prior work that either learned representations from videos for only object and verb recognition [22] or learned motor control in simulation [47], our work demonstrates the applicability of end-to-end learned video representations for real-world robotic control.
@@
We present the key contributions of our work as follows: (1) We present a transformer-based policy to encode video task specification, demonstrated by either robot or human agent embodiments (§II). (2) We encourage alignment between the prompt and robot video representations using three contrastive auxiliary losses during training (§II-E) (3) Through actual robot experiments, we find our video-conditioned policy is better than baselines on human prompt videos. Furthermore, our policy is better at cross-object motion transfer (§III).
@@
Figure 2: Dataset creation. (top row) Here we show a Robot-Robot video pair for placing the rxbar into top drawer. We similarly pair existing robot-robot videos performing the same task. (middle row) Here, we show Hindsight Human-Robot paired videos for picking a Coke can from the bottom drawer and placing it on the counter task. We use the task instructions from robot trajectories, ask human participants to perform the task and record a demonstration video from the robot’s perspective/view. (bottom row) Here, we show a Co-located Human-Robot pair of videos for placing the pipe wrench in the toolkit. We record a human demonstration and a robot teleoperation in the same workspace. We use different workspaces to perform the same task instruction, thus collecting paired videos with visually diverse prompts and robot state observations. More details in Section II-B.
@@ II-A Preliminaries
Our objective is to design a robotic system that takes in a prompt video of a manipulation task and outputs actions that accomplish the task demonstrated in the video. This system must infer the underlying task from the prompt video (which might have a different setup or embodiment than the robot) and then manipulate the objects in its environment to achieve the inferred task. Our model’s inputs are a prompt video $V$ and the robot state $S_{t}=\{x_{i}\}_{i=t-k-1}^{t}$ where $x_{i}$ is the frame from the robot’s camera stream at time $i$, $k$ is the maximum number of historical frames, and $t$ is the current timestep We train a policy $\pi(a_{t}|S_{t},V)$ that infers the underlying task from $V$ and predicts task-relevant action $a_{t}$. We need a dataset of paired prompt videos and robot trajectories to train this model. Below, we will discuss in detail how to create paired datasets.
@@ II-B Datasets
[…]
@@
(1) Robot-Robot: We pair existing robot-robot videos of the same task. For this pairing, we consider two videos to match if they perform the same task in different settings. We define “task” based on natural language instructions for recording robot trajectories. These instructions typically consist of one or two verbs surrounded by nouns, such as ‘place water bottle upright’, ‘move the coke can to the green chip bag’ or ‘open top drawer’. Note that we use language instructions only for pairing and as an input to the robot policy. The objective of this pairing is two-fold: first, to leverage the already labeled and collected dataset of robot trajectories, and second, to ensure robots can imitate the same task but in different environment.
@@
(2) Hindsight Human-Robot: Here, we use the task instructions from the robot trajectories dataset, ask one to five human participants to perform the task, and record a demonstration video from the robot’s perspective/view. The instructions are the same as before, but there is a significant embodiment and speed variability due to different humans performing the task with left or right hands and at a randomized robot camera angle. This provides us with a lot of paired data for training the policy with the available set of instructions in the robot dataset, without having to collect new robot trajectories.
@@
(3) Co-located Human-Robot: In this case, humans and robots perform the same task in the same workspace. We used this approach to collect human demonstrations and robot trajectories in diverse spaces such as a living space with sofas, a meeting room with whiteboards, a hardware workstations with toy tools, a kitchen with countertop, a refrigerator and a sink, a storage supplies area, and more.
@@
We show examples of paired videos of the prompt and robot demo from each of the three datasets in Figure 2. As can be seen, there is a considerable difference in the backgrounds and distractor objects in the Hindsight Human-Robot and Co-located Human-Robot datasets. A different complexity arises when comparing the first approach (Robot-Robot), where the actor is a robot with the same morphology, to the other two cases where the human is the actor in the prompt videos.
Each source represents a different level of difficulty and time to collect. Pairing existing robot datasets requires no extra data collection but does not involve any human demonstrations. Our second data involves asking humans to mimic existing robot trajectories. Hindsight human videos made data collection easier as they do not need robot teleoperation data. However, this did not increase the diversity of tasks in the data set. Lastly, we collect data with humans and robots in the same environment. While collecting co-located paired data is a presumed gold standard, it is very time and labor-intensive compared to the previous two approaches. Thus, it forms a small fraction of our overall training set. After combining all the datasets, we have $\sim$100k robot videos and $\sim$10k human videos covering the tasks introduced in RT-1 [8] and RT-2 [9]. The robot-robot dataset makes up more than 90% of the entire dataset. This dataset is publicly available [35]. We provide a Python Notebook in Supplementary Material for accessible visualization of the paired videos used in training.
@@ II-C Model Architecture
Figure 3: Architecture. Our model takes the prompt video and the robot’s current observations as the input, encodes those into token embeddings for the prompt video and the robot’s state, cross-attends to produce state-prompt encoding, and translates it into the expected robot action at the current timestep. More details in Section II-C).
@@
Our policy takes the prompt video and the current robot state as inputs and outputs robot actions. It consists of four modules: (1) prompt video encoder, (2) robot state encoder, (3) state-prompt encoder, and (4) robot action decoder. The entire architecture is illustrated in Figure 3, and each of the modules is detailed below:
@@
(1) Prompt Video Encoder encodes the video demonstration provided as a reference to convey the desired task semantics. The prompt video encoder implicitly learns to infer what task to perform and how to do it. The prompt encoder consists of a per-frame Image encoder $\phi_{p}$ (ViT [15]) followed by a Perceiver Resampler [1, 20] $\psi_{p}$. The output of the prompt encoder $\psi_{p}(\phi_{p}(V))=z_{prompt}$ is a set of $N$ tokens of d-dimension to condition the policy with the task-relevant attributes from the video.
@@
(2) Robot State Encoder encodes the current state of the robot given the current frame and last $k$ frames as input. Note that this module also encodes information about the objects and environment visible to the robot. The architecture is similar to the prompt encoder: a per-frame Image encoder $\phi_{s}$ followed by a Perceiver Resampler $\psi_{s}$. Identical to the prompt encoder’s outputs, the output of the state encoder is $\psi_{s}(\phi_{s}(S_{t}))=z_{state}$ that encodes the latent environment and robot state information from the history of recent observations. We use the same image encoder weights for both (1) and (2), that is, $\phi_{p}\!=\!\phi_{s}\!=\!\phi$. The role of the image encoder $\phi$ is to capture spatial visual information in each frame. The perceiver resampler enables temporal learning across frames and reduces the number of video tokens passed into the action decoder.
@@
(3) State-Prompt Encoder takes the prompt video encoding $z_{prompt}$ and robot state encoding $z_{state}$ and outputs a task encoding relevant for action prediction, which we call prompt-aware state tokens $z_{state|prompt}$. Here, the state encoding acts as queries, and the prompt video encoding acts as keys and values. Intuitively, the state-prompt encoder enables the fusion of the state and prompt information. Suppose a prompt video shows picking up an apple in the basket, and the current state contains an apple, a banana, and an orange. The State-Prompt Encoder cross-attends and learns which object to attend to in the state based on the prompt video. Capturing interdependencies between prompt and state is crucial for the next step of action decoding.
@@
(4) Robot Action Decoder predicts the action vector $a_{t}$ for the current state $S_{t}$ such that it completes the task shown in the prompt video $V_{p}$. The action decoder is a transformer decoder architecture that uses the fixed action position tokens [53] as input queries and the prompt-aware state tokens $z_{state|prompt}$ for keys and values. The size of the action position embedding is $N\times d$ where $N$ is the number of action dimensions, and $d$ is the transformer embedding dimension. More details on action vector in §II-D.
@@
The action position embeddings cross-attend to the prompt-aware state tokens to predict the target binned action values as output. Each output token of the action decoder corresponds to an action dimension for the mode, arm, and base. We project each token embedding to 256 dimensions, and a softmax layer is applied on the top to obtain the bin corresponding to the target action vector. Unlike prior work [8, 9] that use autoregressive action decoding that requires multiple forward passes during inference, we use action position embeddings for one forward pass prediction like in ACT [53]. Instead of predicting one action for the next timestep, we follow the approach outlined in [22, 53] and train the policy with a prediction horizon of four steps. We always use the action bin with the highest probability, that is, argmax over predicted probabilities, to choose the action value for execution.
[…]